OpenStudy (anonymous):

antiderivative of (1/(2x))-(Sqrt[2*(e^x)])

OpenStudy (turingtest):

\[\int\limits {1\over2x}-\sqrt{2e^x}dx\]right?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

gotta think about it...

OpenStudy (across):

\[\int\frac{1}{2}x^{-1}dx-4\int e^{\frac{1}{2}x}dx\]Does that make it any easier to solve?

OpenStudy (turingtest):

no across I don't think you rewrote it correctly... let me try

OpenStudy (turingtest):

\[\int\limits (2x)^{-1}dx-\sqrt2\int\limits e^xdx\]that should be pretty doable

OpenStudy (across):

You're right, the 4 should be a sqrt(2)

OpenStudy (turingtest):

I'm having breakfast, can you guys try it?

OpenStudy (across):

\[\frac{1}{2}\int x^{-1}-\sqrt{2}\int e^{\frac{1}{2}x}\]

OpenStudy (turingtest):

yeah, that's right across, I messed up earlier too

OpenStudy (across):

\[\frac{1}{2}\ln x-2\sqrt{2}\sqrt{e^x}\]

OpenStudy (turingtest):

add a +C and across is correct